OpenStudy (anonymous):

I'm trying to solve the following inequality and list in inverval notation: 8 - |2x - 1| >= 6

OpenStudy (anonymous):

I first tried dividing all by 8 and originally got [-1/4,1/2] but that answer was wrong.

OpenStudy (texaschic101):

8 - |2x - 1 | >= 6 --- subtract 8 -|2x - 1| >= 6 - 8 -|2x - 1| >= -2 -- divide by -1 |2x - 1| <= 2 (2x - 1) <= 2 2x <= 2 + 1 2x <= 3 x <= 3/2 (2x - 1) >= -2 2x >= -2 + 1 2x >= -1 x >= -1/2 -1/2 <= x <= 3/2 interval notation : [-1/2,3/2]